1.(-4+2i)+(6-3i)
2.6√-12
3.√-8*√-32
4.i^65

5.i^89

Answer by stanbon(48545)

1.(-4+2i)+(6-3i) = 2 -i
--------------------------------
2.6√-12 = 6*2sqrt(3)i
-------------------------
3.√-8*√-32 = 2sqrt(2)i * 4sqrt(2)i = 8*2*(-1) = -16
---------------------------
4.i^65 = i^64 * i = 1*i = i
----------------------------------
5.i^89 = i^88 * i = 1*i = i
===============================
Cheers,
Stan H.